# Better Instruction-Following Through Minimum Bayes Risk

### v2

@@ 1 Introduction
From our MBR distillation experiments, we find that self-training with Direct Preference Optimisation (DPO) (Rafailov et al., 2024) on preference pairs selected using MBR decoding with Prometheus-2-7B substantially improves greedy decoding performance. For instance, MBR self-trained Llama2-13b improves by +7.1% on AlpacaEval 2.0 and +0.90 on MT-Bench relative to its baseline SFT counterpart when evaluated using only greedy decoding, far surpassing the corresponding gains from BoN self-training. We also find that MBR self-trained models evaluated with greedy decoding generally match and sometimes exceed the performance of their base models evaluated with MBR decoding, thereby demonstrating that MBR distillation is an effective way of mitigating the inference-time costs of MBR decoding while retaining improved performance.
@@ 4.1.1 Training an SFT Model
We use 3000 randomly-drawn samples from UltraChat-200k (Ding et al., 2023) for SFT. UltraChat is a highly diverse conversational instruction-following dataset created using GPT-3.5-Turbo. Each sample consists of multi-turn prompts and responses, although we only take the first turn of each sample in order to simplify experimentation. We designate our SFT models as sft. See Appendix I for a summary of our SFT hyperparameters.
@@ 4.1.2 Iterative DPO on MBR-Decoded Outputs
Having obtained SFT models, we now conduct DPO iteratively to improve the models on their own MBR-decoded outputs. We start by randomly drawing a further 3000 prompts from UltraChat (excluding the samples that have already been selected for SFT). Next, we generate $N_{\mathrm{cand}}=12$ candidate outputs from our sft models using these prompts. We then score the candidate outputs using a utility metric and form preference pairs from the highest-scoring and lowest-scoring outputs. This preference pair dataset is then used for DPO on the sft models, yielding dpo-MBR-1. We repeat this process twice more, each time using the latest dpo models as the base model paired with a fresh set of 3000 prompts, yielding the models dpo-MBR-2 and dpo-MBR-3. See Appendix I for a summary of our DPO hyperparameters and Appendix G.4 for experimental results from using another preference pair selection strategy.
@@ 4.2 Results
DPO self-training significantly improves model performance We report the results of our self-training experiment in the left subtable of Table 4. We find that three rounds of MBR self-training with DPO significantly improves model performance, with the 7B dpo-3-MBR model outperforming the 13B sft model with greedy decoding on both AlpacaEval 2.0 and MT-Bench. The improvements saturate by the third round of self-training as measured on MT-Bench (6.11 vs 6.14 (7B) and 6.73 vs 6.75 (13B) in Table 4), although there appears to be room for further improvement on AlpacaEval 2.0 (7.22 vs 8.86 (7B) and 13.9 vs 15.3 (13B) in Table 4). Both dpo-3-MBR models outperform their sft-full counterparts, which suggests that training on MBR-decoded outputs is more beneficial than SFT on a larger split of UltraChat. The dpo-3-MBR models also generally outperform sft with MBR decoding, and this is especially prominent for MT-Bench, which suggests that DPO on MBR-decoded outputs enables models to recover and then exceed their MBR-decoding performances. We find that DPO on BoN-decoded outputs also improves model performance, although less so than DPO with MBR-decoded outputs. We attribute this to the relative strength of MBR decoding.
SFT self-training yields smaller gains than DPO self-training We experiment with iterative SFT self-training, using the 7B sft model. We document our results in the right subtable of Table 4. We use the same sets of prompts as for DPO and select as our SFT labels the highest-scoring sample as determined by MBR. As before, we conduct three rounds of iterative self-training, yielding sft-1-MBR, sft-2-MBR and sft-3-MBR. We find that SFT training yields significantly less improvement than DPO. This indicates that MBR self-training benefits most from preference learning, where the model is trained to contrast its highest- and lowest-consensus quality outputs.
@@ 5 Related Work
MBR decoding MBR decoding has been explored in the context of machine translation using a variety of machine translation metrics such as BLEU, COMET (Rei et al., 2020) and BLEURT (Sellam et al., 2020) as utility metrics, with promising results (Freitag et al., 2022; 2023; Farinhas et al., 2023; Stanojević & Sima’an, 2014; Eikema & Aziz, 2020). Prior works (Bertsch et al., 2023; Jinnai et al., 2023) also study MBR decoding for summarisation, using ROUGE and BERTScore as metrics and BART as the language model. Suzgun et al. (2022) apply MBR decoding to several tasks, including summarisation, machine translation and three different BIG-Bench tasks (Srivastava et al., 2023). None of these works explore the use of MBR decoding in the more open-ended and challenging instruction-following domain, nor do they consider using LLM judges as utility metrics.
LLM Judges Based on the strong instruction-following capabilities of LLMs, recent works explore directly prompting LLMs to judge responses from other LLMs (Li et al., 2023; Zheng et al., 2023). Follow-up works suggest that training on the evaluation traces of stronger models may equip smaller models with strong evaluation capabilities (Kim et al., 2023; 2024; Zhu et al., 2023b; Vu et al., 2024; Wang et al., 2024b). These works mainly focus on training LLMs to produce scoring decisions matching those of humans. In our work, instead of viewing evaluation as an end goal, we explore methods to utilise the evaluation capabilities of LLM judges as supervision to improve the instruction-following capabilities of LLMs.
@@
Self-training Self-training is a promising avenue for model improvement as it enables training without external labelled data. Gulcehre et al. (2023) introduce an algorithm that iteratively generates samples from a policy and then updates the policy using offline RL. Yuan et al. (2024b) train models to score their own outputs, and then use these scores to create preference datasets which they distil back into the model using DPO. Huang et al. (2022) train models on their own highest-confidence outputs as determined by majority voting. Self-training on MBR-decoded outputs has also been explored for machine translation. Finkelstein et al. (2024) train models with SFT on their own MBR and quality estimation outputs for machine translation and demonstrate that this yields improvements over baseline models. Wang et al. (2024a) use MBR to generate targets for sequence-level distillation, again for machine translation. Yang et al. (2024) use DPO to upweight the model’s own MBR generations, allowing them to recover much of their original MBR performances using only greedy decoding.

### v3

@@ 1 Introduction
From our MBR distillation experiments, we find that self-training with Direct Preference Optimisation (DPO) (Rafailov et al., 2024) on preference pairs selected using MBR decoding (Yang et al., 2024) with Prometheus-2-7B substantially improves greedy decoding performance. For instance, MBR self-trained Llama2-13b improves by +7.1% on AlpacaEval 2.0 and +0.90 on MT-Bench relative to its baseline SFT counterpart when evaluated using only greedy decoding, far surpassing the corresponding gains from BoN self-training. We also find that MBR self-trained models evaluated with greedy decoding generally match and sometimes exceed the performance of their base models evaluated with MBR decoding, thereby demonstrating that MBR distillation is an effective way of mitigating the inference-time costs of MBR decoding while retaining improved performance.
@@ 4.1.1 Training an SFT Model
We use 3000 randomly-drawn samples from UltraChat-200k (Ding et al., 2023) for SFT. UltraChat is a highly diverse conversational instruction-following dataset created using GPT-3.5-Turbo. Each sample consists of multi-turn prompts and responses, although we only take the first turn of each sample in order to simplify experimentation. We designate our SFT models as sft.
@@ 4.1.2 Iterative DPO on MBR-Decoded Outputs
Having obtained SFT models, we now conduct DPO to improve the models on their own MBR-decoded outputs, an approach first proposed by Yang et al. (2024) for improving machine translation. We start by randomly drawing a further 3000 prompts from UltraChat (excluding the samples that have already been selected for SFT). Next, we generate $N_{\mathrm{cand}}=12$ candidate outputs from our sft models using these prompts. Following Yang et al. (2024), we then score the candidate outputs using a utility metric and form preference pairs from the highest-scoring and lowest-scoring outputs. This preference pair dataset is then used for DPO on the sft models, yielding dpo-MBR-1. We extend upon Yang et al. (2024) by iteratively repeating this process twice more, each time using the latest dpo models as the base model paired with a fresh set of 3000 prompts, yielding the models dpo-MBR-2 and dpo-MBR-3. See Appendix I for a summary of our SFT and DPO hyperparameters and Appendix G.4 for experimental results from using another preference pair selection strategy.
@@ 4.2 Results
DPO self-training significantly improves model performance We report the results of our self-training experiment in the left subtable of Table 4. We find that three rounds of MBR self-training with DPO significantly improves model performance, with the 7B dpo-3-MBR model outperforming the 13B sft model with greedy decoding on both AlpacaEval 2.0 and MT-Bench. The improvements saturate by the third round of self-training as measured on MT-Bench (6.11 vs. 6.14 (7B) and 6.73 vs. 6.75 (13B) in Table 4), although there appears to be room for further improvement on AlpacaEval 2.0 (7.22 vs. 8.86 (7B) and 13.9 vs. 15.3 (13B) in Table 4). Both dpo-3-MBR models outperform their sft-full counterparts, which suggests that training on MBR-decoded outputs is more beneficial than SFT on a larger split of UltraChat. The dpo-3-MBR models also generally outperform sft with MBR decoding, and this is especially prominent for MT-Bench, which suggests that DPO on MBR-decoded outputs enables models to recover and then exceed their MBR-decoding performances. We find that DPO on BoN-decoded outputs also improves model performance, although less so than DPO with MBR-decoded outputs. We attribute this to the relative strength of MBR decoding.
SFT self-training yields smaller gains than DPO self-training We experiment with iterative SFT self-training, using the 7B sft model. We document our results in the right subtable of Table 4. We use the same sets of prompts as for DPO and select as our SFT labels the highest-scoring sample as determined by MBR, following Finkelstein et al. (2024). As before, we conduct three rounds of iterative training, yielding sft-1-MBR, sft-2-MBR and sft-3-MBR. We find that SFT training yields significantly less improvement than DPO. This indicates that MBR self-training benefits most from preference learning, where the model learns to contrast its highest- and lowest-quality outputs.
@@ 5 Related Work
MBR decoding MBR decoding has been explored in the context of machine translation using a variety of translation metrics such as COMET (Rei et al., 2020) and BLEURT (Sellam et al., 2020), with promising results (Freitag et al., 2022; 2023; Farinhas et al., 2023; Stanojević & Sima’an, 2014). Prior works (Bertsch et al., 2023; Jinnai et al., 2023) also study MBR decoding for summarisation, using ROUGE and BERTScore as metrics. Suzgun et al. (2022) apply MBR decoding to several tasks, including summarisation, machine translation and three different BIG-Bench tasks (Srivastava et al., 2023). None of these works explore the use of MBR decoding in the more open-ended instruction-following domain, nor do they consider using LLM judges as utility metrics.
LLM judges Based on the strong instruction-following capabilities of LLMs, recent works explore prompting LLMs to judge responses from other LLMs (Li et al., 2023; Zheng et al., 2023). Follow-up works suggest that training on the evaluation traces of strong models may equip smaller models with strong evaluation capabilities (Kim et al., 2023; 2024; Zhu et al., 2023b; Vu et al., 2024; Wang et al., 2024b). These works focus on training LLMs to produce scoring decisions matching those of humans. In our work, instead of viewing evaluation as an end goal, we explore utilising the evaluation capabilities of LLM judges as supervision to improve instruction-following LLMs.
@@
Self-training Self-training is a promising avenue for model improvement as it enables training without external labelled data. Gulcehre et al. (2023) introduce an algorithm that iteratively generates samples from a policy and then updates the policy using offline RL. Yuan et al. (2024b) train models to score their own outputs, and then use these scores to create preference datasets which they distil back into the model using DPO. Huang et al. (2022) train models on their own highest-confidence outputs as determined by majority voting. Self-training on MBR-decoded outputs has also been explored for machine translation. Finkelstein et al. (2024) train models with SFT on their own MBR and quality estimation outputs for machine translation and demonstrate that this yields improvements over baseline models. Wang et al. (2024a) use MBR to generate targets for sequence-level distillation, again for machine translation. Yang et al. (2024) are the first to use DPO to upweight the model’s own MBR generations, allowing them to recover much of their original MBR performances on machine translation using only greedy decoding.